If the fifth overtone of a closed pipe is in unison with the fifth overtone of an open pipe, then the ratio of the length of closed pipe to open pipe will be
(a) 11:12
(b) 12:11
(c) 13:11
(d) 11:13

Grade:Upto college level

2 Answers

Arun
3 years ago

3rd overtonr of closed = 2nd harmonic of open

7\left [ \frac{V}{4l_{c}} \right ]= 2\left [ \frac{V}{l_{0}} \right ]

\frac{l_{c}}{l_{0}}= \frac{7}{4}

 

Closed organ pipe -

A closed organ pipe is a cylindrical tube having an air colume with one end closed.

- wherein

Condition of constructive interference

l=\left ( 2n+1 \right )\frac{\lambda }{4}    

Vikas TU
3 years ago
means both have same frequency for closed pipe 
f=11c/4l
for  open pipe
f=6c/2l
'11c/4l
=6c/2l
'l'/l=24/22=12/11Regards